OpenStudy (anonymous):

Solve the following equation: 6y – 20 = 2y – 4.

OpenStudy (anonymous):

6y-20=2y-4 Since 2y contains the variable to solve for, move it to the left-hand side of the equation by subtracting 2y from both sides. 6y-20-2y=-4 Since 6y and -2y are like terms, add -2y to 6y to get 4y. 4y-20=-4 Since -20 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 20 to both sides. 4y=20-4 Subtract 4 from 20 to get 16. 4y=16 Divide each term in the equation by 4. (4y)/(4)=(16)/(4) Simplify the left-hand side of the equation by canceling the common factors. y=(16)/(4) Simplify the right-hand side of the equation by simplifying each term. y=4

OpenStudy (anonymous):

6y-2y = -4+20 4y = 16 y=4